The company is considering a project involving a purchase of new equipment. Change the data area of your worksheet to match the following:

Cost of new equipment needed: $370,000

Working capital needed: $45,000

Overhaul of equipment in four years: $30,000

Salvage value of equipment in five years: $25,000

Annual revenues and costs:

Sales revenues: $430,000

Cost of goods sold: $250,000

Out-of-pocket operating costs: $70,000

Discount rate: 15%

1. What is the net present value of the project?

2. The internal rate of return is between what two whole discount rates?

3. Reset the discount rate to 15%. Suppose the salvage value is uncertain. How large would the salvage value have to be to result in a positive net present value?
Answer #1

1. NPV of the project

Year 0 1 2 3 4 5
Cost of equipment -370000
Working capital -45000
Overhaul of equipment -30000
Salvage Value 25000
Release of working capital 45000
Annual cash inflows 110000 110000 110000 110000 110000
Net cash flows -415000 110000 110000 110000 80000 180000
PV @ 15% 1 0.870 0.756 0.658 0.572 0.497
Present Value -415000 95652 83176 72327 45740 89492
Net Present Value -28613

2. Calculation of IRR

At IRR, NPV = 0

0 = -415000 + 110000/ (1 + IRR) + 110000/ (1 + IRR)^2 + 110000/ (1 + IRR)^3 + 80000/ (1 + IRR)^4 + 180000/ (1 + IRR)^5

IRR = 12.20%

3. Suppose the salvage value is uncertain. How large would the salvage value have to be to result in a positive net present value

Calculation of NPV without salvage value:

Year 0 1 2 3 4 5
Cost of equipment -370000
Working capital -45000
Overhaul of equipment -30000
Release of working capital 45000
Annual cash inflows 110000 110000 110000 110000 110000
Net cash flows -415000 110000 110000 110000 80000 155000
PV @ 15% 1 0.870 0.756 0.658 0.572 0.497
Present Value -415000 95652 83176 72327 45740 77062
Net Present Value -41043

For NPV to be positive the Present Value of salvage should be greater or equal to NPV i.e $41,043

PV of salvage = Salvage Value * PV (15%, 5 year)

41,043 = Salvage Value * 0.497

Salvage Value = 41,043 / 0.497 = $ 82,551